Analyzing Manufacturing Cost Accounts

Clapton Company manufactures custom guitars in a wide variety of styles. The following incomplete ledger accounts refer to transactions that are summarized for May:

Materials
May 1 Balance 36,900 May 31 Requisitions (a)
31 Purchases 148,400


Work in Process
May 1 Balance (b) 31 Completed jobs (f)
31 Materials (c)
31 Direct labor (d)
31 Factory overhead applied (e)


Finished Goods
May 1 Balance 0 May 31 Cost of goods sold (g)
31 Completed jobs (f)


Wages Payable
May 31 Wages incurred 120,000


Factory Overhead
May 1 Balance 22,300 May 31 Factory overhead applied (e)
31 Indirect labor (h)
31 Indirect materials 19,700
31 Other overhead 61,200

In addition, the following information is available:

  1. Materials and direct labor were applied to the following jobs in May:
    Job No. Style Quantity Direct Materials Direct Labor
    101 AF1 200 $24,800 $15,000
    102 AF3 420 42,040 27,000
    103 AF2 170 12,630 7,000
    104 VY1 260 41,340 26,000
    105 VY2 190 27,780 17,000
    106 AF4 120 9,280 4,000
    Total 1,360 $157,870 $96,000
  2. Factory overhead is applied to each job at a rate of 130% of direct labor cost.
  3. The May 1 Work in Process balance consisted of two jobs, as follows:
    Job No. Style Work in Process,
    May 1
    101 AF1 $5,900
    102 AF3 16,400
    Total $22,300
  4. Customer jobs completed and units sold in May were as follows:
    Job No. Style Completed in
    May
    Units Sold
    in May
    101 AF1 X 160
    102 AF3 X 336
    103 AF2 0
    104 VY1 X 218
    105 VY2 X 158
    106 AF4 0

Required:

1. Determine the missing amounts associated with each letter and complete the following table. If required, round amounts to the nearest dollar. If an answer is zero, enter in "0". Enter all amounts as positive numbers.

Job No. Quantity May 1
Work in
Process
Direct
Materials
Direct
Labor
Factory
Overhead
Total Cost Unit Cost Units Sold Cost of Goods Sold
No. 101 $ 5,900 $ 24,800 $ 15,000 $ $ $ $
No. 102 16,400 42,040 27,000
No. 103 12,630 7,000
No. 104 41,340 26,000
No. 105 27,780 17,000
No. 106 9,280 4,000
Total $22,300 $157,870 $96,000 $ $ $

a. Materials Requisitions $

b. Work in Process Beginning Balance $

c. Direct Materials $

d. Direct Labor $

e. Factory overhead applied $

f. Completed jobs $

g. Cost of goods sold $

h. Indirect labor $

2. Determine the May 31 balances for each of the inventory accounts and factory overhead. Use the minus sign to indicate any credit balances.

Materials $
Work in Process $
Finished Goods $
Factory Overhead $
Answer #1
1
a. Material requisitions=Total direct materials+Indirect materials=157870+19700=$ 177570
b. Work in process beginning balance=Cost of Job 101 and 102 as on May 1=5900+16400=$ 22300
c. Direct materials=Total direct materials cost=$ 157870
d. Direct labor=$ 96000
e. Factory overhead applied=Direct labor*130%=96000*130%=$ 124800
f. Completed job=Cost of completed jobs
Job No. 101 102 104 105 Total
Beginning balance 5900 16400
Direct materials 24800 42040 41340 27780
Direct labor a 15000 27000 26000 17000
Factory overhead applied (a*130%) 19500 35100 33800 22100
Total 65200 120540 101140 66880 353760
Completed job=$ 331460
g. Cost of goods sold
Job No. 101 102 104 105 Total
Total cost a 65200 120540 101140 66880
Quantity produced b 200 420 260 190
Cost per unit c=a/b 326 287 389 352
Units sold d 160 336 218 158
Cost of goods sold e=c*d 52160 96432 84802 55616 289010
Cost of goods sold=$ 289010
h. Indirect labor=Wages incurred-Direct labor=120000-96000=$ 24000
Job No. Quantity May 1 work in process Direct materials Direct labor Factory overhead Total cost Unit cost Units sold Cost of goods sold
a b c d e=d*130% f=b+c+d+e g=f/a h i=g*h
No. 101 200 5900 24800 15000 19500 65200 326 160 52160
No. 102 420 16400 42040 27000 35100 120540 287 336 96432
No. 103 170 12630 7000 9100 28730 169 0 0
No. 104 260 41340 26000 33800 101140 389 218 84802
No. 105 190 27780 17000 22100 66880 352 158 55616
No. 106 120 9280 4000 5200 18480 154 0 0
Total 22300 157870 96000 124800 400970 289010
2 May 31 balances
Materials:
$ $
Beginning balance 36900
Purchases 148400
185300
Less:
Direct materials 157870
Indirect materials 19700 177570
May 31 balance 7730
Work in process:
$
Beginning balance 22300
Add:
Direct materials 157870
Direct labor 96000
Factory overhead applied 124800
400970
Less: cost of completed job 331460
May 31 balance 69510